# Methodology, Investigation, Software, Visualization, Writing – original draft

### v1

@@ 4.3.2 Experimental Results
The enhancement results of reclassification are illustrated in Fig.(6), where we observed that ScPace outperformed ScReclassify on the simulated datasets, especially at higher mislabeling rates, such as 40% and 30%. However, we noted that Psupertime performed poorly on original mislabeled datasets compared to the calibrated timestamps obtained using ScPace and ScReclassify. This observation suggests that Psupertime lacks robustness when confronted with extremely noisy datasets.
The enhancement results from the deletion method in ScPace, shown in Fig.( 7), indicate a clear improvement on the simulated datasets on both types of noise. This finding suggests that ScPace effectively removes noisy labeled datasets, leading to enhanced performance in supervised pseudotime analysis. Interesingly, we discovered that Psupertime is sensitive to two types of noise, as the mislabeling rate increases, the correlation values drop significantly.
Moreover, On real time-series ScRNA-seq datasets. Both ScPace and ScReclassify successfully enhanced the performance of supervised pseudotime analysis, however, ScPace demonstrated better performance compared to ScReclassify across all three datasets. This indicates that our proposed method is effective not only in simulated contexts but also in real-world applications, further validating its robustness and utility in supervised pseudotime analysis.
@@
We also conducted experiments utilizing the deletion method. The results showed that the datasets calibrated by ScPace, through the deletion of identified mislabeled cells, significantly enhanced the performance of supervised pseudotime analysis. This improvement indicates that the deletion of these cells was not random; rather, it involved a systematic identification of cells that were likely to contribute noise due to mislabeling. Consequently, the removal of these non-representative cells contributed to a clearer and more accurate reconstruction of pseudotime ordering. Thus, our findings suggest that the effective reclassification and deletion of mislabeled cells within the context of time-series ScRNA-seq analyses can lead to substantial improvements in the accuracy and reliability of pseudotime estimations.
@@
4.3.3 Sensitivity Analysis
4.4 Analysis of mislabeled cells detected By ScPace
4.4.1 Simulated Datasets
After training the original datasets using ScPace, we identified samples with high loss values that were detected as potential mislabeled cells. These samples were subsequently assessed further to determine their validity. The reclassification results, as illustrated in Fig. (8), demonstrate that ScPace effectively detected these potential mislabeled cells, allowing us to create a cleaned dataset by either removing the detected cells or reclassifying them in terms of optimized parameters.
However, it is crucial to carefully tune the hyperparameters while conducting timestamps calibration method to achieve the best results. We observed that if the parameter $C$ is set too low, ScPace tends to remove a larger number of cells than expected, which can lead to the loss of valuable data. Conversely, setting $C$ too high may result in insufficient removal of mislabeled cells, allowing erroneous entries to persist in the dataset. This sensitivity to hyperparameter tuning highlights the need for a balanced approach to parameter selection, ensuring that the calibration process optimally identifies and mitigates mislabeled cells without unnecessarily compromising the dataset’s integrity.
4.4.2 Real time-series ScRNA-seq Datasets
To further investigate whether the deleted cells were randomly removed, we analyzed the detected and normal cells within real time-series ScRNA-seq datasets. In Fig. (9), the red circles highlight potential mislabeled cells (denoted as $v=0$). Our results show that ScPace successfully detected the majority of overlapping cells in the real time-series ScRNA-seq datasets, underscoring its capability to identify label noise effectively.
@@ 4.4.2 Real time-series ScRNA-seq Datasets
We extended our analysis to other datasets besides GSE90047, and observed that the detected cells identified by ScPace were frequently affected by technical issues such as dropout events and data noise. We found no evidence of biological factors that could explain the occurrence of noisy timestamps in these datasets. This underscores the significance of addressing technical artifacts in ScRNA-seq data to ensure the accuracy of biological interpretations.
In conclusion, the presence of label noise in time-series ScRNA-seq data can lead to confusion and potentially degrade the performance of supervised pseudotime analysis. Given the challenges posed by these noisy timestamp labels, we do not recommend simply deleting the detected samples in the GSE90047 dataset. Instead, we advocate for reclassification methods to calibrate the cells into more reliable timestamp labels. This approach is preferable because the observed label noises do not appear to stem from technical issues but rather may reflect biological variability. By employing reclassification techniques, we can enhance the integrity of the analysis while retaining valuable information from the dataset.
@@ 5 Discussion
Previously, researchers have devoted effort to find techniques that better distinguish the proliferation status of cardiomyocytes such as stable, cell-cycle re-entry, true cell division. However, that detection still remains unspecific because of the complexity of cell cycle variants, e.g., binucleation[43]. Future research should concentrate on employing robust timestamp automatic annotation methods to effectively assess the potential proliferative capabilities of cardiomyocytes, which is critical need in the domain of human heart regeneration. Thus, ScPace may serve as a viable alternative for the automatic annotation of cardiomyocyte proliferation status. which could help maintain robustness even in the presence of label noises.
@@ Acknowledgments
This work was partly supported by the National Natural Science Foundation of China (62306051,62481540175), the Fundamental Research Funds for the Central Universities (NJ2022028), and the Scientific and the Technological Research Program of Chongqing Municipal Education Commission (KJQN202300718). Weikai Li was partly supported by the National Natural Science Foundation of China (62076124).

### v2

@@ 4.3.2 Experimental Results
The enhancement results of reclassification are illustrated in Fig.(12), where we observed that ScPace outperformed ScReclassify and CL on the simulated datasets, especially at higher mislabeling rates, such as 40% and 30%. However, we noted that Psupertime performed poorly on original mislabeled datasets compared to the calibrated timestamps obtained using ScPace and ScReclassify. This observation suggests that Psupertime lacks robustness when confronted with extremely noisy datasets.
The enhancement results from the deletion method in ScPace, shown in Fig.(13), indicate a clear improvement on the simulated datasets on both types of noise. Moreover, compared with the enhancement results of CL, ScPace performed better. This finding suggests that ScPace effectively removes noisy labeled datasets, leading to enhanced performance in supervised pseudotime analysis. Interestingly, we discovered that Psupertime is sensitive to two types of noise, as the mislabeling rate increases, the correlation values drop significantly.
Moreover, on real time-series ScRNA-seq datasets. ScPace, ScReclassify and CL successfully enhanced the performance of supervised pseudotime analysis through reclassification, however, ScPace demonstrated better performance compared to ScReclassify and CL across all three datasets. This indicates that our proposed method is effective not only in simulated contexts but also in real-world applications, further validating its robustness and utility in supervised pseudotime analysis.
@@
We also conducted experiments utilizing the deletion method on real time-series ScRNA-seq. The results showed that the datasets calibrated by ScPace, through the deletion of identified mislabeled cells, significantly enhanced the performance of supervised pseudotime analysis compared with CL. This improvement indicates that the deletion of these cells was not random; rather, it involved a systematic identification of cells that were likely to contribute noise due to mislabeling. Consequently, the removal of these non-representative cells contributed to a clearer and more accurate reconstruction of pseudotime ordering. Thus, our findings suggest that the effective reclassification and deletion of mislabeled cells within the context of time-series ScRNA-seq analyses can lead to substantial improvements in the accuracy and reliability of pseudotime estimations.
@@
5 Case Studies
5.1 Identification of different maturation state in hepatoblast differentiation
@@
In Fig. (14), the red circles highlight potential mislabeled cells (denoted as $v=0$). Our results show that ScPace successfully detected the majority of overlapping cells in GSE90047, underscoring its capability to identify label noise effectively.
Given the challenges posed by these noisy timestamp labels, we do not recommend simply deleting the detected samples in the GSE90047 dataset. Instead, we advocate for reclassification methods to calibrate the cells into more reliable timestamp labels. This approach is preferable because the observed label noises do not appear to stem from technical issues but rather may reflect biological variability. By employing reclassification techniques, we can enhance the integrity of the analysis while retaining valuable information from the dataset.
5.2 Illustration of cardiomyocytes developmental status after myocardial infraction
We performed a primary practice in illustrating cardiomyocytes developmental status after myocardial infraction. The datasets of the cardiomyocytes gathered from GEO database under accession number GSE126128[43], GSE131181[44] and GSE132658[45], with 2404 cells and 6 time points(E7.75, E8.25, E9.25, E10.5, E13.5, E16.5) and furthermore preprocessed using Scanpy package. The datasets were then split into training (75%) and validation (25%) sets.
Based on the training results reflected in the confusion matrix shown in Fig.(15), we observed a degree of ambiguity between E8.25 and E9.25. This is likely attributable to the minimal developmental changes occurring during this stage, making it challenging to accurately distinguish between these two time points.
For testing sets, we gathered 2082 cells from National Center for Biotechnology Information Sequence Read Archive with BioProject accession number PRJNA595398[46]. In terms of the prediction results from the testing sets shown in Fig.(16), we found that the concentration of cardiomyocytes after injury was shifting to the earlier period and the peak at embryonic day 13.5 was specifically increased with recovery process. In addition, our results suggested that obstacles in cardiomyocyte dedifferentiation might occur between status embryonic day 10.5 and day 13.5 due to the density showed little change among earlier periods. Consequently, our results verified cardiomyocyte wound go through the incomplete re-programming process in genomic aspect after myocardial infraction and the number of cardiomyocytes at unmatured status wound increase over time.
6 Computational timing
To assess the computational costs of the model on timestamp calibration. The model is trained on a research server with an NVIDIA GPU GeForce RTX 3080 Ti and 12th Gen Intel(R) Core(TM) i7-12700KF CPU with 125Gi . The computational timing of timestamp calibration on both swap mislabeling and random mislabeling are shown in Table.(7) and Table.(8), and the computational timing of real time-series datasets are shown in Table.(6) . The results indicate that both ScPace and CL exhibit significantly longer computational times when dimensionality reduction techniques are not employed, compared to ScReclassify, which incorporates PCA in its initial steps. However, we observed that when dimensionality reduction techniques are applied to ScPace, its computational time becomes even shorter than that of ScReclassify.
@@ 7 Discussion
Another limitation of ScPace arises when the mislabeling rate exceeds 50%. Under such conditions, where more than half of the cells are mislabeled, the model loses its capacity to reliably identify mislabeled samples. Consequently, future research could prioritize the development of more robust methodologies capable of addressing extreme levels of label noise.
@@ Acknowledgments
This work was partly supported by the National Natural Science Foundation of China (62306051,62481540175,82270249), Group Building Scientific Innovation Project for Universities in Chongqing (CXQT21021), Joint Training Base Construction Project for Graduate Students in Chongqing (JDLHPYJD2021016), the Scientific and the Technological Research Program of Chongqing Municipal Education Commission (KJZD-K202400703,KJQN202300718) and the Natural Science Foundation of Chongqing (CSTB2023NSCQ-LZX0092)